Circular No. 7719

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ej IN UGC 3829 A. Dimai, Cortina d’Ampezzo, Italy, reports his discovery of an appar-

ent supernova (mag ∼ 16.5) on CCD images taken on Sept. 17.12 (limiting mag ∼ 18.5) and 18.14 UT (limiting mag ∼ 19.5) with a 0.51-m reflector in the course of the CROSS program (cf. IAUC 7373). SN 2001ej is located at α = 7^{h}23^{m}43^{s}.23, δ = +33^{o}26′38′′.0 (equinox 2000.0), which is ∼ 4′′ west and 7′′ north of the center of UGC 3829. A CROSS image taken on Apr. 5 showed nothing at this position (limiting mag ∼ 18.5), and nothing is visible on Palomar Sky Survey red and blue plates.

SUPERNOVA 2001ek IN ANONYMOUS GALAXY F. R. Boffi, Space Telescope Science Institute (STSI), on behalf of the

Supernovae at High Z Team at the STSI, reports Hubble Space Telescope (HST) observations of the probable supernova reported by P. Challis on IAUC 7668. Observations taken with the HST-WPC2 instrument on Aug. 3 confirm the detection of the supernova and yield photometric measurements of [F555W] = 25.45 ± 0.05, [F675W] = 23.99 ± 0.03, and [F814W] = 23.45 ± 0.03. Boffi adds: “The observed colors, and the amounts and signs of the magnitude changes in 44 days, are best explained in terms of a type-IIP supernova (at redshift 0.25–0.40) that was discovered ∼ 20–40 days past maximum and whose spectrum is affected by little or no reddening. The corresponding rest-frame B absolute magnitude at maximum is estimated to be −17.1 ± 0.6, which is typical of a type-IIP supernova (Patat et al. 1994, A.Ap. 282, 731).”

V4739 SAGITTARII CCD photometry by A. C. Gilmore and P. M. Kilmartin at Mount

John (cf. IAUC 7702): Sept. 2.358 UT, V = 11.55, U–B = −0.71, B–V = −0.32, V–R = +1.67, V–I = +1.14; 8.333, 12.76, −0.74, −0.31, +1.41, +1.02; 9.356 (marginal conditions; clouds), 12.96, −0.71, −0.34, +1.39, +0.70; 10.423, 13.23, −0.78, −0.41, +1.37, +1.06; 11.365, 13.50, −0.82, −0.47, +1.45, +0.9; 12.348, 13.59, −0.80, −0.38, +1.32, +0.6.

V4740 SAGITTARII Visual magnitude estimates: Sept. 12.250 UT, 7.0 (J. Bedient, Hon-

olulu, HI); 14.024, 7.3 (R. Shida, São Paulo, Brazil); 15.92, 7.5 (J. Carvajal, Avila, Spain); 17.625, 7.3 (W. Souza, São Paulo, Brazil).

2001 September 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green